OpenStudy (anonymous):

x+y=-2, 2x-2y=-8 (-3 1) substitution method

OpenStudy (anonymous):

x=-3, y=1. plug it in. -3+1=-2, 2(-3)-2(1)=-8 what are you supposed to do?

OpenStudy (anonymous):

x+y=0 3x-3y=0 (-1,1) my answer is -1=0,-6=0 is that correct ??

OpenStudy (anonymous):

i dont understand what you did..

OpenStudy (anonymous):

ok this is the question x+y=0 3x-3y=0 (-1,1) i have the answer for it can you solve it and i will check if we have the same answer